# A continuum mechanics approach for the deformation of non-Euclidean origami generated by piecewise constant nematic director fields

### v1

@@ Abstract
We merge classical origami concepts with active actuation by designing origami patterns whose panels undergo prescribed metric changes. These metric changes render the system non-Euclidean, inducing non-zero Gaussian curvature at the vertices after actuation. Such patterns can be realized by programming piecewise constant director fields in liquid crystal elastomer (LCE) sheets. In this work, we address the geometric design of both compatible reference director patterns and their corresponding actuated configurations. On the reference configuration, we systematically construct director patterns that satisfy metric compatibility across interfaces. On the actuated configuration, we develop a continuum mechanics framework to analyze the kinematics of non-Euclidean origami. In particular, we fully characterize the deformation spaces of three-fold and four-fold vertices and establish analytical relationships between their deformations and the director patterns. Building on these kinematic insights, we propose two rational designs of large director patterns: one based on a quadrilateral tiling with alternating positive and negative actuated Gaussian curvature, and another combining three-fold and four-fold vertices governed by a folding angle theorem. Remarkably, both designs achieve compatibility in both the reference and actuated states. We anticipate that our geometric framework will facilitate the design of non-Euclidean/active origami structures and broaden their application in active metamaterials, soft actuators, and robotic systems.
@@ 1 Introduction
Shape programming has emerged as a fundamental topic across diverse fields, including robotics, metamaterials, soft electronics, and geometric design choi2019programming ; filipov2015origami . Among various shape programming techniques, origami—the ancient art of paper folding—offers a promising platform for designing complex configurations through coordinated folding lines Levi2016Programming ; lang1996computational ; lang2011origami . Since paper is nearly inextensible, traditional origami involves isometric deformations that preserve distances within each panel. As a result, the surfaces remain developable, with zero Gaussian curvature, and the design space becomes highly constrained. The resulting nonlinearity and over-constrained geometry make many origami-related problems computationally challenging; for instance, determining rigid-foldability with optional creases is known to be NP-hard Akitaya_Demaine_Horiyama_Hull_Ku_Tachi_2020 . Due to these challenges, much research focuses on special classes of origami, such as rigidly and flat-foldable quadrilateral mesh origami feng2020designs , where inverse design frameworks have been developed dang2022inverse . Additional work on rigid foldability includes tachi2009generalization ; lang2018rigidly ; izmestiev2017classification ; hayakawa2024panel , offering accurate predictions for actuation, though typically requiring external control mechanisms ai2024easy .
@@
Combining origami geometry with metric actuation seems to provide a new paradigm, potentially enabling shape-morphing systems that are both predictable, functionally programmable and remotely controllable. A seminal example is provided by Plucinsky et al. plucinsky2016programming ; plucinsky2018patterning (also shown in Fig. 1), where origami panels are augmented with panels patterned by piecewise-constant LCE director fields. Upon actuation, the panels elongate along the director by a factor $\lambda_{\parallel}$ and contract transversely by $\lambda_{\perp}$. Although each panel remains flat, the actuation induces non-isometric deformations, giving rise to non-trivial Gaussian curvature at the vertices o2006elementary ; duffy2020defective . The resulting configuration often bifurcates into distinct “up” and “down” states due to the induced curvature. While the experiments validate theoretical predictions, their design space remains limited due to symmetry constraints, limiting applicability in scenarios requiring more complex actuation. We also find that the systematic study of active non-Euclidean origami is quite limited in the literature. To enable more versatile functionality, it is crucial to develop a generalized design framework for active origami (also generically resulting in non-Euclidean origami waitukaitis2019non ) that can systematically handle director pattern design and accurately predict deformation behavior.
@@
In this paper, we focus on the geometric and kinematic aspects of such a framework. We restrict our attention to a purely geometric regime, assuming that each panel undergoes a uniform, piecewise-constant metric change prescribed by the director field, and remains flat and rigid after actuation. Within this framework, two main challenges arise. First, the reference director pattern must be compatible, meaning the prescribed length changes must match across each crease. A simple case occurs when the crease is the bisector of adjacent directors or orthogonal to it—a condition referred to as metric compatibility feng2020evolving . However, patterning compatible director fields becomes nontrivial when accounting for topological and geometric interactions among multiple creases. Second, the kinematics of the actuated configuration must be predicted accurately. Ideally, the kinematics of the actuated shape can be predicted directly from the reference director pattern. A key question remains: even when the reference director pattern is compatible, does the actuated configuration maintain compatibility?
@@
We address these challenges by first analyzing the fundamental conditions for metric compatibility and then studying the kinematics of actuated three-fold and four-fold intersections. We generalize previous continuum mechanics approaches feng2020designs ; zou2024kinematics to derive closed-form kinematic equations for these non-Euclidean vertices. Specifically, we show that the actuated three-fold vertices are rigid with no free folding angle and are fully determined by the director configuration, while the four-fold vertices have one free folding angle, with the remaining angles following from a unified analytical formula. Using these insights, we investigate multi-node interactions. For example, a pattern with two connected three-fold vertices reveals hidden degrees of freedom in director programming. For larger-scale patterns, we propose a marching algorithm to design quadrilateral director tilings with alternating positive and negative actuated Gaussian curvature. Additionally, we construct another class of compatible patterns composed of both three-fold and four-fold vertices (Fig. 1(b)), governed by a folding angle theorem. Surprisingly, despite the complexity of these systems, the active origami patterns are compatible in both the reference and actuated configurations and offer substantial design flexibility.
@@
This paper is organized as follows. Section 2 introduces the deformation of director patterns and the conditions for metric compatibility. Sections 3 and 4 analyze the kinematics of three-fold and four-fold vertices, respectively. Section 5 presents two design strategies for large-scale director fields and their actuated configurations. Finally, Section 6 concludes the paper and discusses future research directions.
@@ 2 Deformation of constant director fields and metric compatibility condition
which implies ${\bf U}_{{\bf n}}{\bf n}=\lambda_{\parallel}{\bf n}$ and ${\bf U}_{{\bf n}}{\bf n}_{\perp}=\lambda_{\perp}{\bf n}_{\perp}$. Notice that ${\bf U}_{{\bf n}}$ is invariant under the transformation ${\bf n}\rightarrow-{\bf n}$, which means the director ${\bf n}$ is actually double-arrowed, and changing ${\bf n}$ to $-{\bf n}$ results in the same pattern and actuation.
@@
Next, we consider the metric-compatible interface with tangent ${\bf t}$ between the two constant director fields ${\bf n}_{1}$ and ${\bf n}_{2}$ satisfying ${\bf n}_{1}\nparallel{\bf n}_{2}$. The interface will serve as the crease of origami in the following sections. If we assume that the interfaces are compatible before and after actuation (the lengths are identical from the two sides), one can obtain the metric-compatibility condition

### v2

@@ Abstract
We merge classical origami concepts with active actuation by designing origami patterns whose panels undergo prescribed metric changes. These metric changes render the system non-Euclidean, inducing non-zero Gaussian curvature at the vertices after actuation. Such patterns can be realized by programming piecewise constant director fields in liquid crystal elastomer (LCE) sheets. In this work, we address the geometric design of both compatible reference director patterns and their corresponding actuated configurations. On the reference configuration, we systematically construct director patterns that satisfy metric compatibility across interfaces. We prove the existence and uniqueness of compatible director fields at a vertex for the generic case, up to orthogonal duals. The Gaussian curvature of the actuated vertex is computed based on the compatible director fields. On the actuated configuration, we develop a continuum mechanics framework to analyze the kinematics of non-Euclidean origami. In particular, we fully characterize the deformation spaces of three-fold and four-fold vertices and establish analytical relationships between their deformations and the director patterns. Building on these kinematic insights, we propose rational designs of large director patterns: one based on a quadrilateral tiling with alternating positive and negative actuated Gaussian curvature, and the other combining three-fold and four-fold vertices governed by a folding angle theorem. Remarkably, both designs achieve compatibility in both the reference and actuated states. We also propose a design strategy for active metamaterials based on the periodic non-Euclidean origami. The active metamaterials can have two modes of motions by folding or stimulating. We anticipate that our geometric framework will facilitate the design of non-Euclidean/active origami structures and broaden their application in active metamaterials, soft actuators, and robotic systems.
@@ 1 Introduction
Shape programming has emerged as a fundamental topic across diverse fields, including robotics, metamaterials, soft electronics, and geometric design choi2019programming ; filipov2015origami ; dang2025kirigami . Among various shape programming techniques, origami—the ancient art of paper folding—offers a promising platform for designing complex configurations through coordinated folding lines Levi2016Programming ; lang1996computational ; lang2011origami ; misseroni2024origami . Since paper is nearly inextensible, traditional origami involves isometric deformations that preserve distances within each panel. As a result, the surfaces remain developable, with zero Gaussian curvature, and the design space becomes highly constrained. The resulting nonlinearity and over-constrained geometry make many origami-related problems computationally challenging; for instance, determining rigid-foldability with optional creases is known to be NP-hard Akitaya_Demaine_Horiyama_Hull_Ku_Tachi_2020 . Due to these challenges, much research focuses on special classes of origami, such as rigidly and flat-foldable quadrilateral mesh origami feng2020designs , where inverse design frameworks have been developed dang2022inverse . Additional work on rigid foldability includes tachi2009generalization ; lang2018rigidly ; izmestiev2017classification ; hayakawa2024panel , offering accurate predictions for actuation, though typically requiring external control mechanisms ai2024easy .
@@
Combining origami geometry with metric actuation seems to provide a new paradigm, potentially enabling shape-morphing systems that are both predictable, functionally programmable and remotely controllable. In this work, the in-plane non-isometric deformation induces non-zero Gaussian curvature at the vertices in the deformed configuration, while the out-of-plane deformation is geometrically restricted by folding. A seminal example is provided by Plucinsky et al. plucinsky2016programming ; plucinsky2018patterning (also shown in Fig. 1), where origami panels are augmented with panels patterned by piecewise-constant LCE director fields. Upon actuation, the panels elongate along the director by a factor $\lambda_{\parallel}$ and contract transversely by $\lambda_{\perp}$. Although each panel remains flat, the actuation induces non-isometric deformations, giving rise to non-trivial Gaussian curvature at the vertices o2006elementary ; duffy2020defective . The resulting configuration often bifurcates into distinct “up” and “down” states due to the induced curvature. While the experiments validate theoretical predictions, their design space remains limited due to symmetry constraints, limiting applicability in scenarios requiring more complex actuation. We also find that the systematic study of active non-Euclidean origami is quite limited in the literature. To enable more versatile functionality, it is crucial to develop a generalized design framework for active origami (also generically resulting in non-Euclidean origami waitukaitis2019non ) that can systematically handle director pattern design and accurately predict deformation behavior.
@@
In this paper, we focus on the geometric and kinematic aspects of such a framework. We restrict our attention to a purely geometric regime, assuming that each panel undergoes a uniform, piecewise-constant metric change prescribed by the director field, and remains flat and rigid after actuation. Within this framework, two main challenges arise. First, the reference director pattern must be compatible, meaning the prescribed length changes must match across each crease. A simple case occurs when the crease is the bisector of adjacent directors or orthogonal to it—a condition referred to as metric compatibility feng2020evolving . However, patterning compatible director fields becomes nontrivial when accounting for topological and geometric interactions among multiple creases. For example, it is not obvious that a compatible director fields at a vertex exists. Second, the kinematics of the actuated configuration must be predicted accurately. Ideally, the kinematics of the actuated shape can be predicted directly from the reference director pattern. A key question remains: even when the reference director pattern is compatible, does the actuated configuration maintain compatibility?
@@
We address these challenges by first analyzing the fundamental conditions for metric compatibility, proving the existence of the compatible director fields at a general $n$-fold vertex, and then studying the kinematics of actuated three-fold and four-fold intersections. By satisfying the metric compatibility condition, the length changes of the crease are identical from both sides. Under this condition, we prove that compatible director fields always exist at a $n$-fold vertex. They are characterized as the generic case with two determined director fields which are orthogonal to each other (orthogonal dulas) and the degenerate case with a continuous distribution of the director fields. We then generalize previous continuum mechanics approaches feng2020designs ; zou2024kinematics to derive closed-form kinematic equations for these non-Euclidean vertices. Specifically, we show that the actuated three-fold vertices are rigid with no free folding angle and are fully determined by the director configuration, while the four-fold vertices have one free folding angle, with the remaining angles following from a unified analytical formula. Using these insights, we investigate multi-node interactions. For example, a pattern with two connected three-fold vertices reveals hidden degrees of freedom in director programming. For larger-scale patterns, we propose a marching algorithm to design quadrilateral director tilings with alternating positive and negative actuated Gaussian curvature. Additionally, we construct another class of compatible patterns composed of both three-fold and four-fold vertices (Fig. 1(b)), governed by a folding angle theorem. Surprisingly, despite the complexity of these systems, the active origami patterns are compatible in both the reference and actuated configurations and offer substantial design flexibility. Moreover, building on the periodic quadrilateral pattern, we propose a design for active metamaterials that can deform through folding and external stimulation. These active metamaterials have the potential to be coarse-grained to have a new shell/plate theory mcinerney2025coarse ; li2025nonlinear ; XU2024105832 .
@@
This paper is organized as follows. Section 2 introduces the deformation of director patterns and the conditions for metric compatibility. Section 3 studies the patterning of compatible director fields at a general $n$-fold vertex. Sections 4 and 5 analyze the kinematics of three-fold and four-fold vertices, respectively. Section 6 presents two design strategies for large-scale director fields and their actuated configurations. A design strategy for active metamaterials is also proposed. Finally, Section 7 concludes the paper and discusses future research directions.
@@ 2 Deformation of constant director fields and metric compatibility condition
which implies ${\bf U}_{{\bf n}}{\bf n}=\lambda_{\parallel}{\bf n}$ and ${\bf U}_{{\bf n}}{\bf n}_{\perp}=\lambda_{\perp}{\bf n}_{\perp}$. Notice that ${\bf U}_{{\bf n}}$ is invariant under the transformation ${\bf n}\rightarrow-{\bf n}$, which means the director $\bf n$} is actually double-arrowed, and changing $\bf n$} to $\bf n$} results in the same pattern and actuation.
@@
Next, we consider the metric-compatible interface with tangent $\bf t$} between the two constant director fields ${\bf n}_{1}$ and ${\bf n}_{2}$ satisfying ${\bf n}_{1}\nparallel{\bf n}_{2}$. The interface will serve as the crease of origami in the following sections. If we assume that the interfaces are compatible before and after actuation (the lengths are identical from the two sides) and $\lambda_{\parallel,\perp}$ are uniform throughout the entire pattern, one can obtain the metric-compatibility condition
@@
3 Patterning a compatible $k$-fold director vertex
We first consider a general $k$-fold intersection in 2D, where $k$ distinct director fields ${\bf n}_{1},\dots,{\bf n}_{k}$ (${\bf n}_{i}\nparallel{\bf n}_{j}$ if $i\neq j$) are separated by $k$ creases ${\bf t}_{1},\dots,{\bf t}_{k}$. At each crease, the metric compatibility is satisfied as $|{\bf U}_{{\bf n}_{i}}{\bf t}_{i}|=|{\bf U}_{{\bf n}_{i-1}}{\bf t}_{i}|$. In the following, we provide two patterning strategies, prescribing either the director fields or the creases. Then we compute the concentrated Gaussian curvature after actuation based on the 2D director pattern.
3.1 Two patterning strategies
Suppose that $k$ director fields ${\bf n}_{1},\dots,{\bf n}_{k}$ (${\bf n}_{i}\nparallel{\bf n}_{j}$ if $i\neq j$) are prescribed. We design creases ${\bf t}_{i}$ between ${\bf n}_{i}$ and ${\bf n}_{i-1}$ which satisfies the metric compatibility. As discussed above, such ${\bf t}_{i}$ can be written as Eq. (5), meaning that ${\bf t}_{i}$ can bisect the two adjacent directors or be perpendicular to the bisector. However, the branches $\sigma$ and $\tilde{\sigma}$ cannot be arbitrarily chosen. As shown in Fig. 3, the creases must satisfy the topological condition
$\operatorname{sign}[({\bf t}_{i}\times{\bf t}_{i+1})\cdot{\bf e}_{3}]=+1,\quad{\text{for }i=1,2,...,k}$ (6)
to preserve the counterclockwise sequence of ${\bf t}_{i}$. The topological condition itself is subtle to be satisfied, as the branches $\sigma$ and $\tilde{\sigma}$ can not be explicitly implied by this condition. We have to substitute specific $\sigma$ and $\tilde{\sigma}$ and check the condition. Thus, a more realistic strategy is to prescribe the creases and then distribute the director fields.
Now suppose that $k$ creases are given by ${\bf t}_{i+1}={\bf R}_{{\bf e}_{3}}(\alpha_{1}+\dots+\alpha_{i}){\bf e}_{1}$ and ${\bf t}_{1}={\bf e}_{1}$, where $\alpha_{i}$ satisfies the topological condition $0<\alpha_{i}<\pi$ and $\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}=2\pi$. We first assume that a compatible director field ${\bf n}_{i},i=1,2,\dots,k$ exists. An orthogonal dual to this director field is defined as a new field by replacing ${\bf n}_{i}$ with ${\bf n}_{i}^{\perp}={\bf R}_{{\bf e}_{3}}(\pi/2){\bf n}_{i}$ in the corresponding domains, as shown in Fig. 3(a). The following Lemma implies that a compatible director field’s orthogonal dual is also compatible.
Lemma 3.1.
Suppose a $k$-fold director field is compatible at each crease ($|{\bf t}_{i}\cdot{\bf n}_{i-1}|=|{\bf t}_{i}\cdot{\bf n}_{i}|$). Then the orthogonal dual is also compatible ($|{\bf t}_{i}\cdot{\bf R}_{{\bf e}_{3}}(\pi/2){\bf n}_{i-1}|=|{\bf t}_{i}\cdot{\bf R}_{{\bf e}_{3}}(\pi/2){\bf n}_{i}|$).
Proof.
The proof is straightforward by projecting ${\bf n}_{i},{\bf n}_{i-1}$ onto ${\bf t}_{i}$ and ${\bf t}_{i}^{\perp}$ as ${\bf n}_{i-1}=a_{1}{\bf t}_{i}+b_{1}{\bf t}_{i}^{\perp},{\bf n}_{i}=a_{2}{\bf t}_{i}+b_{2}{\bf t}_{i}^{\perp}$. The compatibility $|{\bf t}_{i}\cdot{\bf n}_{i-1}|=|{\bf t}_{i}\cdot{\bf n}_{i}|$ yields $|a_{1}|=|a_{2}|$. Since ${\bf t}_{i}$ and ${\bf t}_{i-1}$ are both unit vectors, we then have $|b_{1}|=|b_{2}|$, which is the compatibility $|{\bf t}_{i}\cdot{\bf R}_{{\bf e}_{3}}(\pi/2){\bf n}_{i-1}|=|{\bf t}_{i}\cdot{\bf R}_{{\bf e}_{3}}(\pi/2){\bf n}_{i}|$ for the orthogonal dual. ∎
Figure 3: (a) An n-fold director field and its orthogonal dual for the generic case. The solution for ${\bf n}_{i}$ is unique up to orthogonal duals. (b) A degenerate case that has a continuous family of solutions for ${\bf n}_{i}$.
The following theorem implies the existence and uniqueness of the compatible director fields for a given set of creases (generic case), up to orthogonal duals.
Theorem 3.1.
Suppose $k$ creases are prescribed by ${\bf t}_{1}={\bf e}_{1}$ and ${\bf t}_{i}={\bf R}_{{\bf e}_{3}}(\alpha_{1}+\alpha_{2}+\dots+\alpha_{i-1}){\bf e}_{1},i=2,\dots,k$, where the sector angles satisfy $0<\alpha_{i}<\pi$ and $\alpha_{1}+\dots+\alpha_{k}=2\pi$. Then generically there are two and only two compatible director fields, which are the orthogonal duals to each other, except for the degenerate cases.
Proof.
Without loss of generality, we have ${\bf t}_{1}={\bf e}_{1}$. Let ${\bf n}_{1}={\bf R}_{{\bf e}_{3}}(\theta){\bf e}_{1},\theta\in(0,\pi)$. To satisfy the metric compatibility $|{\bf t}_{i}\cdot{\bf n}_{i-1}|=|{\bf t}_{i}\cdot{\bf n}_{i}|$, we systematically pattern the directors as ${\bf n}_{i+1}={\bf n}_{i}-2({\bf n}_{i}\cdot{\bf t}_{i+1}){\bf t}_{i+1}$ for $i=1,\dots,k-1$. This will make the compatibility condition ${\bf t}_{i}\cdot{\bf n}_{i-1}=-{\bf t}_{i}\cdot{\bf n}_{i}$ to be satisfied at all creases except ${\bf t}_{1}$. Thus we may check $error=({\bf n}_{k}\cdot{\bf t}_{1})^{2}-({\bf n}_{1}\cdot{\bf t}_{1})^{2}$ for the last compatibility at ${\bf t}_{1}$. Notice that ${\bf n}_{1}$ is a function of $\theta$ and ${\bf t}_{i}$ are given. Thus, the other directors can be derived sequentially, resulting in a function $error(\theta)$. Substituting $\theta$ and $\alpha_{i}$, a direct calculation yields
$error(\theta)=\begin{cases}-\sin[2(\alpha_{1}+\alpha_{3}+\dots+\alpha_{k-1})]\sin[2(\alpha_{1}+\alpha_{3}+\dots+\alpha_{k-1}+\theta)]\quad\text{for even k}\\ -\sin[2(\alpha_{2}+\alpha_{4}+\dots+\alpha_{k-1})]\sin[2(\alpha_{2}+\alpha_{4}+\dots+\alpha_{k-1}+\theta)]\quad\text{for odd k}\end{cases}.$ (7)
Notice that $\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}=2\pi$. Thus $error(\theta)\equiv 0$ if the sum of the even indexed terms or the odd indexed terms in $\alpha_{i}$ is $\pi/2,\pi$ or $3\pi/2$, and $\theta$ can be any value in $(0,\pi)$ (degenerate case in Fig. 3(b)). Otherwise, $error(\theta)=0$ has two solutions for $\theta\in(0,\pi)$. Meanwhile, we need to check ${\bf n}_{k}\times{\bf n}_{1}\neq 0$ to make sure ${\bf n}_{k}$ is not parallel to ${\bf n}_{1}$, and ${\bf t}_{1}$ is a crease. Again, by direct calculation, we have
${\bf n}_{k}\times{\bf n}_{1}=\begin{cases}(0,0,-\sin[2(\alpha_{1}+\alpha_{3}+\dots+\alpha_{k-1}+\theta)])\quad\text{for even k}\\ (0,0,-\sin[2(\alpha_{2}+\alpha_{4}+\dots+\alpha_{k-1})])\quad\text{for odd k}\end{cases}.$ (8)
Taking these two conditions together, we conclude that
  1. 1.
    Generic case: $\alpha_{2}+\alpha_{4}+\dots\neq\pi/2,\pi$ or $3\pi/2$. If $k$ is odd, there are two compatible patterns for ${\bf n}_{i}$ (by satisfying $\sin[2(\alpha_{2}+\alpha_{4}+\dots+\alpha_{k-1}+\theta)]=0$) which are orthogonal duals to each other, since $error(\theta)$ and $error(\theta\pm\pi/2)$ vanish simultaneously (Fig. 3(a)). If $k$ is even, we eventually get ${\bf n}_{k}\parallel{\bf n}_{1}$ by patterning a compatible director filed, which means that ${\bf t}_{1}$ is not a crease and the even $k$ case degenerates into the odd $k$ case.
  2. 2.
    Degenerate case: $\alpha_{2}+\alpha_{4}+\dots=\pi/2,\pi$ or $3\pi/2$. If the sum of the even indexed (or equivalently odd indexed) sector angles is $\pi/2,\pi$ or $3\pi/2$, say $\alpha_{2}+\alpha_{4}+\dots=\pi/2,\pi$ or $3\pi/2$, and $k$ is even, we have $\theta\in(0,\pi)$ is always a solution since $\sin[2(\alpha_{1}+\alpha_{3}+\dots+\alpha_{k-1})]=0$. The number of compatible patterns ${\bf n}_{i}$ is infinite. If $k$ is odd and $\alpha_{2}+\alpha_{4}+\dots=\pi/2,\pi$ or $3\pi/2$, we have ${\bf n}_{k}\times{\bf n}_{1}=0$ (${\bf n}_{k}\parallel{\bf n}_{1}$) and the case is degenerated into the even $k$ case.
The results can be conveniently confirmed by using an alternative Householder transformation (matrix) approach in mirror reflections. ∎
3.2 Concentrated Gaussian curvature at the vertex
In this section, we study the Gaussian curvature concentrated at the vertex after deformation for a compatible director field. As discussed above, we only need to consider two cases: the generic case with odd $k$ and the degenerate case with even $k$. The former case has two solutions for compatible director patterns which are orthogonal duals, and the latter case possesses a continuous variation of director patterns.
For the generic case with odd $k$, we have two solutions for $error(\theta)=0$ which are $\theta=\theta_{0}:=-\alpha_{2}-\alpha_{4}-\dots-\alpha_{k-1}$ and $\theta=\theta_{0}+\pi/2$. Then we have ${\bf n}_{1}={\bf R}_{{\bf e}_{3}}(\theta){\bf e}_{1}$ and other ${\bf n}_{i}$ can be computed successively by ${\bf n}_{i+1}={\bf n}_{i}-2({\bf n}_{i}\cdot{\bf t}_{i+1}){\bf t}_{i+1}$. Since the director we study is piecewise constant, the actuated pattern possesses non-zero Gaussian curvature at the tip and zero Gaussian curvature elsewhere. According to previous work feng2020evolving , the concentrated Gaussian curvature at the tip is $G=2\pi-\sum_{i=1}^{k}\tilde{\alpha}_{i}$, where $\tilde{\alpha}_{i}$ is the actuated sector angle given by Eq. (3). In the following, we use $\beta_{i}=\sum_{j=1}^{i}\alpha_{j}$ for convenience, so that the creases can be written as ${\bf t}_{i}={\bf R}_{{\bf e}_{3}}(\beta_{i}){\bf e}_{1}$. Substituting the creases ${\bf t}_{i}$ and the directors ${\bf n}_{i}$ with $\theta_{0}=-\alpha_{2}-\alpha_{4}-\dots-\alpha_{k-1}$, we have the actuated Gaussian curvature at the tip as a function of $\alpha_{i}$:
$G$ $=$ $G_{odd}(k,\alpha_{1},\dots,\alpha_{k})$
$:=$ $2\pi-\sum_{i=1}^{k}\arccos\frac{(r^{2}+1)\cos\alpha_{i}+(r^{2}-1)\cos\tilde{\beta}_{i}}{\sqrt{[(r^{2}+1)+(r^{2}-1)\cos(\alpha_{i}-\tilde{\beta}_{i})][(r^{2}+1)+(r^{2}-1)\cos(\alpha_{i}+\tilde{\beta}_{i})]}}$
$\text{where }\tilde{\beta}_{i}=2\sum_{j=1}^{i-2}(-1)^{j+1}\beta_{j}+(-1)^{i}\beta_{i-1}+(-1)^{i}\beta_{i}+2\sum_{j=i+1}^{k-1}(-1)^{j}\beta_{j},~\beta_{i}=\sum_{j=1}^{i}\alpha_{j}\text{ and }r=\frac{\lambda_{\parallel}}{\lambda_{\perp}}.$
Figure 4: The Gaussian curvature of orthogonal duals (red and blue) for the degenerate case with (a) $k=6,r=0.9$, $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})=2\pi(0.2,0.13,0.07,0.2,0.23)$ and (b) $k=6,r=2/3$, $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})=2\pi(0.2,0.13,0.07,0.2,0.23)$. The Gaussian curvature of the degenerate case can be modulated within a wide range (positive or negative).
Here if the index $j\notin\{1,2,\dots,k\}$, we set $\alpha_{j}=0$. Given the creases ($\alpha_{i}$) for the generic case, Eq. (3.2) provides an explicit formula for the Gaussian curvature of the tip upon actuation. The Gaussian curvature of the orthogonal dual can be computed by simply replacing $r$ with $1/r$. Since a concentric circle pattern and a radial director pattern generate positive and negative Gaussian curvature respectively for $r<1$ feng2022interfacial , it is natural to ask whether the two orthogonal duals always generate positive and negative Gaussian curvature. Though not explicitly seen from (3.2), the answer is “No”. A numerical test with $k=5,r=1.1$ and $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=2\pi(0.2,0.1,0.05,0.55)$ gives $G=0.445$ and $G=0.807$ for the two orthogonal dulas, which are both positive. Another example is the degenerate case in Fig. 4(a). We refer the reader to duffy2020defective for the Gaussian curvature of a generic director defect.
For the degenerate case with even $k$, the solution $error(\theta)=0$ is a continuous range $\theta\in(0,\pi)$. Thus, we may expect that the actuated Gaussian curvature can also be modulated within a range. Suppose $k$ is even, $\alpha_{2}+\alpha_{4}+\dots=\pi/2,\pi$ or $3\pi/2$ and ${\bf n}_{1}={\bf R}_{{\bf e}_{3}}(\theta){\bf e}_{1}$ for $\theta\in(0,\pi)$. We may follow exactly the same process to compute a Gaussian curvature function $G=G_{even}(k,\alpha_{1},\dots,\alpha_{k})$ for even $k$. We omit the specific derivation and claim that $G_{even}(k,\alpha_{1},\dots,\alpha_{k})$ can be obtained from $G_{odd}(k+1,\alpha_{1},\dots,\alpha_{k},\alpha_{k+1})$ by adding an arbitrary $\alpha_{k+1}$, replacing $\tilde{\beta}_{i}$ with $\tilde{\beta}_{i}+2\theta_{0}-2\theta$ (now $\theta_{0}=-\alpha_{2}-\alpha_{4}-\dots-\alpha_{k}$) and changing the upper limit of summation from $k+1$ to $k$ in $G_{odd}(k+1,\alpha_{1},\dots,\alpha_{k},\alpha_{k+1})$ . The formula has been numerically validated. Figure 4 shows that the Gaussian curvature of the degenerate case can be modulated within a wide range (positive or negative) for the orthogonal duals.